OpenStudy (anonymous):

A skier starting from rest skis straight down a slope 50 meters long in 5.0 seconds. What is the magnitude of the acceleration. ok so i know the answer is " 4 m/s^2 but can anyone show me how to get this answer showing work?

OpenStudy (ash2326):

Skier starts from rest so its inital velocity is 0 u=0 m/s The distance it covers is 50 m S=50 m Time it takes 5 seconds T=5 We have the formula \[S=ut+\frac 12 at^2\] Let's plugin the values of S , u and t \[50=0+\frac 12 a\times(5)^2\] or \[\frac 1 2 a \times 5^2=50\] Can you solve for a or acceleration from here?

OpenStudy (anonymous):

oohhh i get it now thankss !